GIVEN THE FOLLOWING FOR A COMPLETED JOB TOTAL PRIME COSTS TOTAL CONVERSION COSTS OVERHEAD IS 200% OF DIRECT LABOR $40,000 $30,000 COMPUTE DIRECT MATERIAL COST DIRECT LABOR COST OVERHEAD COST TOTAL MANUFACTURING COSTS #2 GIVEN THE FOLLOWING DIRCT MATERIAL= $400,000 DIRECT LABOR = $200,000 OVERHEAD =$100,000 TOTAL MANUFACTURING COST ASSUME ALL WIP WAS COMPLETED AND TRANSFERRED TO FINISHED GOODS MAKE THE ENTRY COST OF GOODS MANUFACTURED ASSUME ALL FINISHED GOODS WERE SOLD MAKE THE ENTRY FOR COST OF GOODS SOLD
Answer #1

ANS 1

Converstion costs includes Direct Labour and Overheads

Let Direct labour be x

Therefore Overheads = 2x

Hence

x + 2x = 30,000

3x = 30,000

x = 10,000

Direct Labour = x =10,000

Overheads = 2x = 2*10,000 = 20,000

Prime Cost includes Direct material and Direct Labour

Hence

Direct Material + Direct Labour = 40,000

Direct Material + 10,000 = 40,000

Direct Material = 40,000-10,000

Direct Material = 30,000

ANS 2

Direct Material = $ 400,000
Direct Labour =$ 200,000
Overheads = $ 100,000
Manufacturing Cost = DM + DL + O/H
                                         = $ 400,000 + $ 200,000 + $ 100,000
                                         = $ 700,000
Entry for transfer of work in progress to finished goods
Finished Goods A/c                                                                         ..Dr $700,000
                       TO Work in Progress A/c $700,000
Entry for transfer of the cost of goods sold to Sales
Cost of Goods Sold A/c                                                                   ..Dr $
Profit A/c                                                                                             ..Dr $
                       TO Sales A/c $